OpenStudy (devonhoward15):

Which counter example disproves the following statement? If a line has a y-intercept of 0, then it has a negative slope.

OpenStudy (devonhoward15):

@Vocaloid

Vocaloid (vocaloid):

to disprove the statement, the function has to have a positive or 0 slope AND a y-intercept of 0 any ideas?

OpenStudy (devonhoward15):

y=2x-1?

Vocaloid (vocaloid):

what's the y-intercept of y = 2x - 1?

OpenStudy (devonhoward15):

y

Vocaloid (vocaloid):

not quite

OpenStudy (devonhoward15):

2x-1

Vocaloid (vocaloid):

remember, the y-intercept is the number y equals when x = 0

Vocaloid (vocaloid):

when x = 0, what is y?

OpenStudy (devonhoward15):

0

Vocaloid (vocaloid):

you sure?

Vocaloid (vocaloid):

y = 2x - 1 if x = 0, what is 2x - 1?

OpenStudy (devonhoward15):

2x-1

Vocaloid (vocaloid):

i mean, what number does that equal?

Vocaloid (vocaloid):

2x - 1 = ___?

OpenStudy (devonhoward15):

1x

Vocaloid (vocaloid):

no, we're looking for a number

Vocaloid (vocaloid):

with no letter

OpenStudy (devonhoward15):

so 1?

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

if x = 0, what is 2x?

OpenStudy (devonhoward15):

0

Vocaloid (vocaloid):

good now what is 2x - 1?

OpenStudy (devonhoward15):

-1

Vocaloid (vocaloid):

great. now, -1 does not equal 0 so we try another answer

OpenStudy (devonhoward15):

ok

Vocaloid (vocaloid):

any other guesses?

OpenStudy (devonhoward15):

y=2x+1

OpenStudy (devonhoward15):

@Vocaloid

Vocaloid (vocaloid):

what's the y-intercept?

Vocaloid (vocaloid):

what number does y equal when x = 0?

OpenStudy (devonhoward15):

3x

Vocaloid (vocaloid):

remember what we said before

Vocaloid (vocaloid):

x = 0 means replace "x" with "0"

Vocaloid (vocaloid):

y = 2x + 1 = 2(0) + 1 = ?

OpenStudy (devonhoward15):

yes

OpenStudy (devonhoward15):

it just doesnt have an x

Vocaloid (vocaloid):

so what do you think the answer is? (hint: it's not 3)

OpenStudy (devonhoward15):

3x?

Vocaloid (vocaloid):

not 3x either

OpenStudy (devonhoward15):

or does that count as 3?

Vocaloid (vocaloid):

when x = 0, what is 2x + 1?

Vocaloid (vocaloid):

hint: if x = 0, what is 2x?

OpenStudy (devonhoward15):

0

Vocaloid (vocaloid):

good so if 2x = 0, what is 2x + 1?

OpenStudy (devonhoward15):

1

Vocaloid (vocaloid):

good, so the y=intercept is 1, so this isn't the answer

Vocaloid (vocaloid):

any other guesses?

OpenStudy (devonhoward15):

the answer is -2x

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

remember the first thing I said

OpenStudy (devonhoward15):

2x than

Vocaloid (vocaloid):

good